OpenStudy (anonymous):

get inverse of x^2 + 4 for x greater than or equal to 0 and verify that f(f^-1 ( x )) = x and f^-1 (f(x)) = x

myininaya (myininaya):

f(x)=x^2+4 doesn't have an inverse function but it does have an inverse relation.

myininaya (myininaya):

Is there a domain restriction on f?

OpenStudy (anonymous):

I have updated the question.

myininaya (myininaya):

 this is a one-to-one function and therefore its inverse is a function and is therefore called a inverse function

myininaya (myininaya):

we have y=x^2+4 do you know hot solve for x?

myininaya (myininaya):

how not hot

OpenStudy (anonymous):

I was able to figure it out! Thanks for the help though

myininaya (myininaya):

oops and i graphed y=x^2-4 but whatever

myininaya (myininaya):

ok cool

OpenStudy (anonymous):

Was going to say haha :p

myininaya (myininaya):

or y=x^2-a where a is positive and x>=0

myininaya (myininaya):

since i don't really label what the y-intercept was

myininaya (myininaya):

didn't*